ROTATIONAL SPECTRUM AND INTERNAL DYNAMICS OF THE Ne-H_{2}S VAN DER WAALS COMPLEX

\underline{YAQIAN LIU} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of several isotopomers of the Ne-H_{2}S van der Waals dimer were measured in the frequency range from 4 to 22 GHz, using a pulsed molecular beam Fourier transform microwave spectrometer. Two K =0 progressions were recorded for the symmetrical isotopomers (with H_{2}S/D_{2}S). This doubling is attributed to an internal rotation motion of the H_{2}S moiety within the complex. The two states can be correlated to the 0_{0}_{0} and 1_{0}_{1} internal rotor states of free H_{2}S and D_{2}S. Only one K =0 progression was measured for Ne-DSH. The excited internal rotor state is no longer metastable since the symmetry constraints no longer apply. An anomalous isotope effect observed in Ar-H_{2}S, where the substitution of hydrogen by deuterium causes an increase in the ground state B rotational constant,^{a} was not observed here, in agreement with the ab initio study by Dykstra and co-workers.^{b} Nuclear quadrupole hyperfine structure was resolved or partially resolved for isotopomers containing ^{33}S and D, respectively, and the corresponding quadrupole coupling constants were determined. These were used to derive information about the intermolecular dynamics. The results are compared with those of Ar-H_{2}S and Ar-H_{2}O.

aH. S. Gutowsky, T. Emilsson, and E. Arunan, J . Chem. Phys. 106, 5309 (1997).

bG. de Oliveria and C. Dykstra, J . Chem. Phys. 110, 289 (1999).